OpenStudy (anonymous):

Part 1: Change 2x + 5y = –30 to slope-intercept form. You must show all work. (2 points) Part 2: State the slope and y-intercept (2 points)

OpenStudy (anonymous):

sllope intercept form is y = mx + b

OpenStudy (anonymous):

isolate y

OpenStudy (anonymous):

5y=-2x-30 Divide by 5 throughout

OpenStudy (anonymous):

I'm sorry, but i figured it out lol

OpenStudy (anonymous):

but here is this one that I totally don't get and never have got it lol

OpenStudy (anonymous):

Solve C = 4e – 7g for g.

OpenStudy (anonymous):

u just have to isolate g

OpenStudy (anonymous):

so first bring 4e to the other side C - 4e = -7g now divide throughout by -7

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

That's too funny!! That's exactly what i just started to write on my paper

OpenStudy (anonymous):

ok now?

OpenStudy (anonymous):

c - 4e/ -7 = g

OpenStudy (anonymous):

yes